# EarthGPT-X: Enabling MLLMs to Flexibly and Comprehensively Understand Multi-Source Remote Sensing Imagery

### v1

@@ III Methodology
The goal is to unify the referring comprehension at any granularity and pixel grounding into a visual prompt framework in the multi-source RS scenarios. To this end, EarthGPT-X is developed to explore a more general and comprehensive visual prompting learning paradigm. The entire EarthGPT-X framework is detailed in this section. Specifically, we first elaborate on the multi-modal content integration mechanism (Section III-A). Subsequently, the cross-domain one-stage fusion training strategy is presented (Section III-B). In addition, the pixel perception module is introduced in Section III-C.
@@ III-A Multi-modal Content Integration
Figure 2: Overview of the proposed EarthGPT-X model architecture.
The multi-modal input stream of the EarthGPT-X is characterized by its tripartite structure, comprising an image global and local perception module, a spatial-aware module, and a text instruction processing module. Those multi-modal content encoding and integration are detailed as follows.
@@ III-A1 Visual Global and Local Perception
As shown in Fig. 2 (a), to enhance visual comprehension, the high-resolution inputted images are divided into sub-images for local visual perception. Concurrently, the images are downsampled into lower resolutions and combined with the sub-images for subsequent mixed encoding. The Mixed Visual Encoders (MVE) are designed to encode inputted images and their variants $I^{i}$ into visual representations $V_{\mathrm{img}}$. To leverage the advantages of different network architectures (CNN and Vision Transformer (ViT) [54]) and integrate the local information and contextual dependencies, the semantically complementary backbones CLIP-ConvNeXt [55] and ViT self-supervised by DINOv2[56] are employed as MVE. This visual encoding process can be formulated as
@@ III-A2 Free-form Spatial-aware
Free-form visual prompts are expected to be flexibly used to guide the model in generating appropriate responses. The shapes of visual prompts include points, boxes, circles, scribbles, etc. To meet the need for flexible interaction at any granularity, a noise augmentation method is developed. Specifically, Gaussian distribution based on the dimensions of the box prompts is sampled to simulate the free-formed shapes. The introduced noise encompasses only a portion of the ground truth or expands it. Subsequently, considering that the visual prompts exhibit varying degrees of sparsity across different shapes, as illustrated in Fig. 2 (b), a randomly initialized ViT is employed as a spatial-aware encoder (SAE) specifically for feature extraction, without any pre-trained weights suitable for image features. The various types of visual prompts $V_{\mathrm{p}}$ are encoded using the designed SAE. The implementation equation can be expressed as follows
@@ III-A3 Text Instructions Tokenizer
As displayed in Fig. 2 (c), To facilitate linguistic comprehension, the text instructions $T_{\mathrm{orig}}$ are encoded by the tokenizer[57] module, and projected into language embeddings $L_{\mathrm{instruct}}$, preparing for the subsequent LLM to capture the nuanced interdependence. The detailed process after disassembly is first to use the subword tokenization[57] method to divide text instructions into subwords, enabling understanding of the structure of the linguistic signals. Subsequently, the Word2Vec[58] is adopted to map each subword to a one-dimensional vector embedding. The tokenizer process can be simply described as
@@ III-A4 Hybrid Signals Mutual Understanding
To bridge the gap and enhance the interplay between dense visual representations and sparse visual prompts features, a hybrid signal mutual understanding mechanism is proposed. Specifically, as displayed in Fig. 2 (d), the image representations $V_{\mathrm{img}}$ and visual prompts features $E_{\mathrm{prompt}}$ are processed by a self-attention layer and feed-forward layer to obtain $V_{\mathrm{sa}}$ and $E_{\mathrm{ff}}$, respectively. The subsequent cross-attention and second self-attention blocks are used to compute the correlation $H_{\mathrm{ca}}$ between hybrid referring information and multi-scale holistic signals. In essence, this designed mechanism enhances detailed and contextual visual perception, leading to a deeper comprehension of the referring areas in holistic images. The whole procedure can be mathematically represented as
@@
where $[V\text{-}P]_{\mathrm{hybrid}}$ is the output of this designed module, containing rich visual information and spatial-aware knowledge from the free-form visual prompts. Then, the hybrid visual features are projected to the same dimension as text embeddings through the vision-to-language modality-align projection layer $\mathcal{F}_{\mathrm{vl\text{-}p}}$. This process can be formulated as
$[V\text{-}P]_{\mathrm{proj}}=\mathcal{F}_{\mathrm{vl\text{-}p}}~{}([V\text{-}P]_{\mathrm{hybrid}}).$ (8)
where $[V\text{-}P]_{\mathrm{proj}}$ involve both multi-scale visual signals and free-form spatial-aware information. Notably, the cross-attention layer is updatable during the training.
@@ III-B Cross-domain One-Stage Fusion Training
The previous training pipeline of MLLM routinely uses a step-wise training strategy[8, 22]. However, in the previous pipeline, extensive manual intervention is required. For example, the researcher needs to manually set the different trainable parameters and datasets across various stages. To alleviate the need for manual efforts without sacrificing performance, the cross-domain one-stage fusion training[11] strategy is designed. Most importantly, this strategy further facilitates multi-domain (i.e., natural, optical, SAR, and infrared) data mutual understanding.
In this stage, the projected hybrid tokens are concatenated with the text instruction embeddings $L_{\mathrm{instruct}}$ to develop the multi-modal LLM input $\mathcal{X}$. Subsequently, the integrated multi-modal input is injected into LLM for alignment and deeply interactive understanding (see Fig. 2 (e)). The process can be expressed as
@@
Unlike other models that introduce additional learnable parameters, to simplify the model structure, the Transformer’s attention layers are unfrozen during the training. Specifically, the self-attention head consists of key $K$, query $Q$, and the value $V$, which are implemented by linear layers. The implementations can be expressed as follows
@@
To summarize, the essence of this process lies in levering multi-domain knowledge to generalize versatile visual reasoning ability into RS multi-source complex scenarios interpretation. Furthermore, through tuning on unified abundant visual prompt-based instruction data, the proposed EarthGPT-X can accomplish referring comprehension of RS imagery with the user’s free-form visual prompts and text instruction-following.
III-C Pixel Perception
To unify referring and grounding in one framework and achieve pixel-level grounding, the pixel encoder-decoder architecture is employed to render the referring object masks, which are shown in Fig. 2 (f). In particular, the special token $pix$ is adopted to represent the pixel grounding masks as embeddings[59]. The embeddings $L_{\mathrm{pix}}$ are generated from the LLM output corresponding to the $pix$ token. Then, the language-to-pixel projection layer $\mathcal{R}_{\mathrm{lp\text{-}p}}$ transforms the embeddings $L_{\mathrm{pix}}$ into the pixel decoder’s feature space. In addition, the grounding encoder is denoted as $\mathcal{G}$, the pixel decoder $\mathcal{P}$. The binary segmentation mask $M$ generated process can be represented as follows
@@ III-C Pixel Perception
In summary, based on the designed tripartite architecture and all-in-one training strategy, the proposed EarthGPT-X can be expected to effectively and comprehensively realize cross-modal and multi-source knowledge comprehension. Moreover, EarthGPT-X can handle multi-grained RS visual tasks, and flexibly unifies referring and grounding in one framework.
@@ IV Dataset Construction
The goal is to endow the proposed model with flexible and comprehension interpretation abilities on multi-source imagery. However, existing infrared and SAR datasets are limited to object detection tasks, hindering the development of various advanced reasoning capabilities. To address this limitation and unleash the potential of MLLMs to comprehend diverse RS data, a multi-source visual prompt-based instruction dataset named M-RSVP is constructed. Particularly, the natural and optical RS image data, which involves a wide range of visual tasks, are utilized for LLM mixed tuning. In essence, multi-domain data fusion facilitates generalizing the universal domain reasoning capabilities into limited SAR and infrared scenarios. The specific M-RSVP construction pipeline is detailed as follows.
IV-A Raw Data Selection
@@
In order to develop multi-source image understanding abilities of the proposed EarthGPT-X, it is indispensable to utilize data from different scenes and sensors, including outdoor natural images, RS optical images, SAR, and infrared images, covering different visual task types. Thus, the natural scene dataset we selected is OpenImages[60] which meets this characteristic and is ideal for training the model’s visual perception capabilities in general environments. Moreover, we also choose Visual Genome[61] for training, to practice reasoning over multi-region and multi-object interactions. Meanwhile, for the RS optical dataset, we use RSVP[22] which is the first RS visual prompting dataset, allowing the model to develop region/point interpretation skills in diverse geographic and environmental settings. Furthermore, SAR detection datasets includes SSDD[62], AIR-SARShip-2.0 [63], SARDet [64], MSAR [65], and SAR-Aircraft [66]. Infrared object detection contains Sea-shipping, Infrared-security [67], Aerial-mancar [68], Double-light-vehicle [69], and HIT-UAV [70]. Leveraging multi-domain datasets is beneficial for equipping the proposed model to obtain robust generalization capabilities and to handle a wide range of visual tasks. In addition, multi-source data integration ensures comprehensive and flexible image understanding across multiple domains and sensor types.
IV-B Multi-source Data Unification
[…]
In conclusion, all the various datasets from different domains are combined for mixed training to deepen mutual understanding. The multiple free-form visual prompts are allowed to instruct the model to generate suitable answers, ensuring the flexibility of human-AI interaction. Furthermore, multi-domain data fusion is beneficial to realize multi-source multi-task interpretation by leveraging the diverse features and scenarios presented in supplementary image types.
@@ V-A1 Implementation Details
The one-stage all-in-one training is conducted in this paper. We use transformer-based decoder-only Llama 2-13B[16] as the basic LLM. The pre-trained DINOv2-ViT L/14 [56] and CLIP-ConvNeXt [55] are adopted as hybrid image encoder MVE. The training image resolution is $448\times 448$. Subsequently, the randomly initialized ViT model is seated as the spatial-aware encoder to refine sparse features. Then, we employ a pre-trained SAM encoder and decoder[20] as the grounding encoder and pixel decoder. The image resolution required to be reshaped for the segmentation model is $1024\times 1024$. Furthermore, we utilize the simple MLPs as the projection layer to realize multi-modal content dimension alignment.
Throughout the training, to keep training efficiency and preserve the generalization ability of backbone models, we only unfreeze the self-attention layers of LLM and the added cross-attention module for tuning. Other modules including visual encoders, projection layers, etc., are kept frozen. The AdamW optimizer[72] with a weight decay of 0.01 and a learning rate of 2e-5 is utilized. The total training is conducted on 8 NVIDIA A100 GPUs with 80 GB memory.

### v2

@@ III Methodology
The EarthGPT-X framework is detailed in this section. We first elaborate on the multi-modal content integration mechanism (Section III-A). Then, the cross-domain one-stage fusion training strategy is presented (Section III-B). The extended pixel perception module is introduced in Section III-C.
@@ III-A Multi-modal Content Integration
Figure 2: Overview of the proposed EarthGPT-X architecture. It integrates images, visual prompts, and text instructions through a hybrid signal interaction mechanism, and employs a unified one-stage fusion training strategy to enable multi-source knowledge integration and multi-level multi-task reasoning.
The multi-modal input stream of the EarthGPT-X is characterized by its tripartite structure, comprising an image global and local perception module, a spatial-aware module, and a text instruction processing module. The multi-modal content encoding and integration are detailed as follows.
@@ III-A1 Visual Global and Local Perception
As shown at the bottom left of Fig. 2, to enhance visual comprehension, the high-resolution input images are divided into sub-images for local visual perception. Concurrently, the images are downsampled into lower resolutions and combined with the sub-images for subsequent mixed encoding. The Mixed Visual Encoders (MVE) are designed to encode input images and their variants $I^{i}$ into visual representations $V_{\mathrm{img}}$. To leverage the advantages of different network architectures (CNN and Vision Transformer (ViT) [53]) and integrate the local information and contextual dependencies, the semantically complementary backbones CLIP-ConvNeXt [54] and ViT self-supervised by DINOv2[55] are employed as MVE. This visual encoding process can be formulated as
@@ III-A2 Free-form Spatial-aware
Free-form visual prompts are expected to be flexibly used to guide the model in generating appropriate responses. The shapes of visual prompts include points, boxes, circles, scribbles, etc. To meet the need for flexible interaction at any granularity, a noise augmentation method is developed. Specifically, Gaussian distribution based on the dimensions of the box prompts is sampled to simulate the free-formed shapes. The introduced noise encompasses only a portion of the ground truth or expands it. Subsequently, considering that the visual prompts exhibit varying degrees of sparsity across different shapes, as illustrated in the bottom right of Fig. 2, a randomly initialized ViT is employed as a spatial-aware encoder (SAE) specifically for feature extraction, without any pre-trained weights suitable for image features. The various types of visual prompts $V_{\mathrm{p}}$ are encoded using the designed SAE. The implementation equation can be expressed as follows
@@ III-A3 Text Instructions Tokenizer
As displayed in left of Fig. 2, to facilitate linguistic comprehension, the text instructions $T_{\mathrm{orig}}$ are encoded by the tokenizer[56] module, and projected into language embeddings $L_{\mathrm{instruct}}$, preparing for the subsequent LLM to capture the nuanced interdependence. The detailed process after disassembly is first to use the subword tokenization[56] method to divide text instructions into subwords, enabling understanding of the structure of the linguistic signals. Subsequently, the Word2Vec[57] is adopted to map each subword to a one-dimensional vector embedding. The tokenizer process can be simply described as
@@ III-A4 Hybrid Signals Mutual Understanding
To bridge the gap and enhance the interplay between dense visual representations and sparse visual prompts features, a hybrid signal mutual understanding mechanism is proposed. Specifically, as displayed in the middle of Fig. 2, the image representations $V_{\mathrm{img}}$ and visual prompts features $E_{\mathrm{prompt}}$ are processed by a self-attention layer and feed-forward layer to obtain $V_{\mathrm{sa}}$ and $E_{\mathrm{ff}}$, respectively. The subsequent cross-attention and second self-attention blocks are used to compute the correlation $H_{\mathrm{ca}}$ between hybrid referring information and multi-scale holistic signals. In essence, this designed mechanism enhances detailed and contextual visual perception, leading to a deeper comprehension of the referring areas in holistic images. The whole procedure can be mathematically represented as
@@
where $V\text{-}P]_{\mathrm{hybrid}}$ is the output of this designed module, containing rich visual information and spatial-aware knowledge from the free-form visual prompts. Then, the hybrid visual features are projected to the same dimension as text embeddings through the vision-to-language modality-align projection layer $\mathcal{F}_{\mathrm{vl\text{-}p}}$. This process can be formulated as
$V\text{-}P]_{\mathrm{proj}}=\mathcal{F}_{\mathrm{vl\text{-}p}}~([V\text{-}P]_{\mathrm{hybrid}}).$ (8)
where $V\text{-}P]_{\mathrm{proj}}$ involve both multi-scale visual signals and free-form spatial-aware information. Notably, the cross-attention layer is updatable during training.
@@ III-B Cross-domain One-Stage Fusion Training
The conventional training pipeline of MLLMs typically follows a two-stage approach [8, 15], which requires extensive manual intervention. In particular, researchers must carefully configure trainable parameters and dataset combinations across different stages, making the process complex and labor-intensive. Different from the prior models, we adopt a cross-domain one-stage fusion training strategy [31], which streamlines the process by uniformly treating all collected datasets while preserving strong performance. More importantly, this unified paradigm facilitates mutual understanding across heterogeneous domains, including natural, optical, SAR, and infrared imagery.
In this stage, the projected hybrid tokens are concatenated with the text instruction embeddings $L_{\mathrm{instruct}}$ to develop the multi-modal LLM input $\mathcal{X}$. Subsequently, the integrated multi-modal input is injected into LLM for alignment and deeply interactive understanding (See Fig. 2 LLM part). The process can be expressed as
@@
Unlike other models that introduce additional learnable parameters, to simplify the model structure, the Transformer’s attention layers are unfrozen during the training. Specifically, the self-attention head consists of key $K$, query $Q$, and value $V$, which are implemented by linear layers. The implementations can be expressed as follows
@@
To summarize, EarthGPT-X leverages multi-domain knowledge to generalize versatile visual reasoning into multi-source RS scenarios. By tuning on abundant unified dual-prompting data, it achieves flexible referring comprehension with free-form visual prompts and text instructions.
III-C Integrated with Pixel Perception Module
To further enhance spatial perception, we integrate a pixel perception module into EarthGPT-X, enabling pixel-level grounding under flexible visual prompts. In particular, the special token pixpix is adopted to represent the pixel grounding masks as embeddings[58]. The embeddings $L_{\mathrm{pix}}$ are generated from the LLM output corresponding to the pixpix token. Then, the language-to-pixel projection layer $\mathcal{R}_{\mathrm{lp\text{-}p}}$ transforms the embeddings $L_{\mathrm{pix}}$ into the pixel decoder’s feature space. In addition, the grounding encoder is denoted as $\mathcal{G}$, the pixel decoder $\mathcal{P}$. The binary segmentation mask $M$ generated process can be represented as follows
@@ III-C Integrated with Pixel Perception Module
In summary, EarthGPT-X, built on a tripartite architecture and an all-in-one training strategy, enables effective cross-modal and multi-source knowledge comprehension. It further supports multi-grained RS visual tasks with flexible integration of referring and grounding in a unified framework.
TABLE II: Statistics of Training Data.
Tasks Raw Data #Samples Domain
Image Captioning RSICD [59] 24.3k Optical
UCM-Captions [60] 10k
Scene Classification RSSCN7 [61] 0.56k
EuroSAT [62] 5.4k
ROC (point) SIOR [63] 200k
Region Captioning DIOR-RSVG [2] 31.5k
ROC (box) AIR-SARShip [64] 1.43k SAR
SSDD [65] 1.86k
SARDet [66] 106k
SAR Aircraft [67] 1.9k
SAR-Ship [68] 33.7k
MSAR [69] 32.5k
ROC (box) Sea-shipping[70] 8.4k Infrared
Infrared-security [71] 9.6k
Aerial-mancar [72] 31.6k
Double-light-vehicle [73] 5.1k
Captioning,
Relationship Analysis
(GPT-assisted)
AIR-SARShip [64] 0.5k SAR
SARDet [66] 106k
SSDD [65] 1.0k
SAR Aircraft [67] 1.9k
Infrared-security [71] 9.6k Infrared
Aerial-mancar [72] 31.6k
Notes: ‘ROC’ denotes Referring Object Classification. The ‘box’ and ‘point’ refer to box-level and point-level visual prompts, respectively.
Figure 3: (a) Box prompts: An example of converting SAR object detection data into multi-modal conversation training data. The training data consists of three parts: the image, the visual prompt, and the text (question and answer). The image is taken directly from the object detection dataset, while the visual prompts are from the ground-truth bounding boxes. The text prompts can be selected from fixed prompt templates (*note: enhanced by GPT-4V). (b) Point prompts: each image in the semantic segmentation dataset is divided into 32 × 32 patches, randomly sampled within each patch as the point prompts, with the category retrieved from the corresponding segmentation map. (c) Free-form prompts: Simulated by applying random noise-based augmentations to regions of interest. These free-form cues emulate flexible user interactions, enabling the model to generalize beyond structured visual prompts.
Figure 4: Illustration of the standardized prompt templates designed for GPT-4V annotation generation. The (a) part shows the public templates for generic data construction, while the (b) part demonstrates the customized task templates, $\mathrm{Role}>$ and $\mathrm{Format}>$, tailored to multi-source RS imagery. These templates embed object categories explicitly, ensuring the correctness and consistency of generated annotations.
@@ IV Dataset Construction
To support comprehensive multi-source RS understanding, we construct a large-scale dataset called M-RSVP, which comprises over 650k samples spanning various tasks, modalities, and annotation formats. The detailed introduction is as follows.
IV-A Multi-Source Multi-Domain Data Construction
@@
For developing multi-source and versatile image understanding abilities of EarthGPT-X, it is indispensable to utilize data from different scenes and sensors, including outdoor natural images, RS optical images, SAR, and infrared images, covering different visual task types. Thus, as shown in Tab. II, we construct a multi-domain collection of raw data covering diverse vision-language tasks. For the RS optical modality, we adopt RSVP [15], the first visual prompting dataset in the RS domain, which enables the model to learn region- and point-level understanding in complex geographic environments. Meanwhile, we incorporate more optical datasets [63, 2] to enrich the diversity of tasks and enhance the model’s generalization in RS scenarios. For SAR and infrared modalities, visual prompting data are mainly derived from publicly available object detection datasets, which are reorganized and annotated in a unified instruction format. Additionally, the natural scene dataset we selected is OpenImages [74] and Visual Genome [75], which meets this characteristic and is ideal for training the model’s visual perception capabilities in general environments. The multi-domain data integration ensures comprehensive and flexible image understanding across multiple domains and sensor types.
IV-B Multi-level Prompting Data Annotation
A meticulously designed multi-strategy annotation method is employed to label diverse multi-domain data, which is uniformly reorganized and formatted into image–visual prompt-text triples. Notably, visual prompts include bounding boxes, points, and free-form types. The annotation process is detailed as follows.
IV-B1 Box prompts
The regional data are based on detection datasets, and ground truth bounding boxes are used as visual prompts to recognize object-level categories. As shown in Fig. 3 (a), an example of a SAR image is provided to illustrate the process. The traditional object detection data are transformed into multi-modal conversational samples, each comprising an image, visual prompts, and text (question-answer pairs). The image is directly sourced from the detection dataset, while the visual prompt is generated by applying slight perturbations to the ground-truth boxes, and the textual prompts are drawn from predefined templates. Specifically, image-level tasks use the boxes $0,0,\mathrm{width},\mathrm{height}]$, which can cover the entire image as the visual prompt.
IV-B2 Point prompts
To enable point-based interaction during training, we construct point prompt annotations by randomly sampling pixels within each ground-truth object mask from semantic segmentation datasets. As illustrated in Fig. 3 (b), each image is divided into $32\times 32$ patches, and the points are randomly sampled within each patch as the visual prompts, where each point corresponds to a pixel-level label annotation.
IV-B3 Free-form Prompts
To enable the model to handle diverse and flexible visual inputs at inference time, we simulate free-form prompts through noise-based augmentation methods. Specifically, we consider both box and point prompts and augment them to reflect the imprecision and looseness of user-defined inputs, as illustrated in Fig. 3 (c). First, for box prompt augmentation, to mimic diverse user-drawn free-form inputs, we apply Gaussian noise to the original ground-truth bounding boxes. Let $\mathcal{B}=(x,y,w,h)$ denote a bounding box parameterized by its top-left corner and dimensions. We generate an augmented box $\mathcal{B}^{\prime}$ via:
$\mathcal{B}^{\prime}=\mathcal{B}+\mathcal{N}(0,\Sigma),\quad\text{where }\Sigma\propto\text{diag}(w^{2},h^{2},w^{2},h^{2}).$ (14)
This noise injection perturbs both the location and scale of the box in a size-aware manner, leading to boxes that may loosely cover, shift from, or extend beyond the target object, thus statistically approximating free-form visual prompts. For point prompts, we simulate user interactions by randomly sampling multiple pixels within the ground-truth object mask and associating them with the same instance label. Given a binary mask $\mathcal{M}\subset\Omega$ representing a target object, we uniformly select $K$ points from the interior of $\mathcal{M}$. These sparse prompts guide the model to localize the target object via point-based reference. Additionally, at inference (see bottom part of Fig. 3 (c)), free-form inputs are pre-processed into bounding boxes, enabling flexible, intuitive user-drawn prompts.
IV-B4 Text Prompts
To support instruction tuning on multi-source RS data, we provide a set of carefully designed natural language instructions that correspond to different visual prompts (e.g., box or point). Each instruction guides the model to perform a specific visual understanding task, such as classification, captioning, or relational reasoning. Designed task instructions are as follows:
  • Please identify the object category of each marked region in the image.
  • Please identify the labels of each marked point in the image.
  • Please provide a brief caption of each marked region in the image.
  • Please provide a detailed caption of each marked point in the image.
  • Please analyze the relationship between all marked regions in the image.
  • Please provide a summarized caption based on all the marked regions in the image.
  • Please analyze how the marked objects interact with each other in the given scene.
In addition, users can employ the <<Mark n>> or <<Region n>> identifiers to specify pixels or regions of interest and freely ask questions about them. These task-specific textual instructions guide the model to perform multiple visual tasks, promoting a more fine-grained understanding of visual content under different granularities.
IV-C GPT4V-assisted Data Generation
To enable high-quality annotations with GPT-4V [76], we followed the design principles of [77] and developed a set of rigorous and standardized prompt templates, as illustrated in Fig. 4. These templates are tailored to distinct tasks: brief captioning, detailed captioning, and inter-relationship analysis. Notably, category names and ground-truth bounding boxes are explicitly embedded in the prompts, ensuring the correctness of the generated annotations by guiding GPT-4V to accurately associate text with visual regions. Beyond correctness, the structured templates also guarantee annotation consistency across large-scale data, avoiding ambiguity and hallucination. To further strengthen the coupling between visual prompts and textual outputs, we adopt Set-of-Marks (SoM) prompting [78], which highlights multiple target objects simultaneously and reinforces spatial-textual grounding. This design enables GPT-4V to produce coherent and complete descriptions even in cluttered scenes. In conclusion, the integration of diverse multi-domain datasets through this rigorous pipeline allows joint training of EarthGPT-X on standardized supervision, enhancing its robustness across RS modalities and improving its generalization to complex real-world scenarios. Additionally, to further ensure reliability, we conduct RS expert inspection on sampled annotations, validating semantic accuracy, coherence, and completeness. This human-in-the-loop verification step complements automatic generation and safeguards the overall quality of the constructed dataset.
@@ V-A1 Implementation Details
The one-stage all-in-one training is conducted in this paper. We use transformer-based decoder-only Llama 2-13B[79] as the basic LLM. The pre-trained DINOv2-ViT L/14 [55] and CLIP-ConvNeXt [54] are adopted as a hybrid image encoder MVE. The training image resolution is $448\times 448$. Subsequently, the randomly initialized ViT model is seated as the spatial-aware encoder to refine sparse features. Then, we employ a pre-trained SAM encoder and decoder[12] as the grounding encoder and pixel decoder. The image resolution required to be reshaped for the segmentation model is $1024\times 1024$. Furthermore, we utilize the simple MLPs as the projection layer to realize multi-modal content dimension alignment.
Throughout the training, to keep training efficiency and preserve the generalization ability of backbone models, we only unfreeze the self-attention layers of LLM and the added cross-attention module for tuning. Other modules, including visual encoders, projection layers, etc., are kept frozen. The AdamW optimizer[80] with a weight decay of 0.01 and a learning rate of 2e-5 is utilized. The total training is conducted on 8 NVIDIA A100 GPUs with 80 GB memory, and the training time is around 310 hours.